# On the Lipschitz Continuity of Set Aggregation Functions and Neural Networks for Sets

### v2

@@ Abstract
The Lipschitz constant of a neural network is connected to several important properties of the network such as its robustness and generalization. It is thus useful in many settings to estimate the Lipschitz constant of a model. Prior work has focused mainly on estimating the Lipschitz constant of multi-layer perceptrons and convolutional neural networks. Here we focus on data modeled as sets or multisets of vectors and on neural networks that can handle such data. These models typically apply some permutation invariant aggregation function, such as the sum, mean or max operator, to the input multisets to produce a single vector for each input sample. In this paper, we investigate whether these aggregation functions are Lipschitz continuous with respect to three distance functions for unordered multisets, and we compute their Lipschitz constants. In the general case, we find that each aggregation function is Lipschitz continuous with respect to only one of the three distance functions. Then, we build on these results to derive upper bounds on the Lipschitz constant of neural networks that can process multisets of vectors, while we also study their stability to perturbations and generalization under distribution shifts. To empirically verify our theoretical analysis, we conduct a series of experiments on datasets from different domains.
@@ 1 Introduction
In the past decade, deep neural networks have been applied with great success to several problems in different machine learning domains ranging from computer vision krizhevsky2012imagenet ; he2016deep to natural language processing vaswani2017attention ; peters2018deep . Owing to their recent success, these models are now ubiquitous in machine learning applications. However, deep neural networks can be very sensitive to their input. Indeed, it is well-known that if some specially designed small perturbation is applied to an image, it can cause a neural network model to make a false prediction even though the perturbed image looks “normal” to humans szegedy2014intriguing ; goodfellow2014explaining .
@@
A key metric for quantifying the robustness of neural networks to small perturbations is the Lipschitz constant. Training neural networks with bounded Lipschitz constant has been considered a promising direction for producing models robust to adversarial examples tsuzuku2018lipschitz ; anil2019sorting ; trockman2021orthogonalizing . However, if Lipschitz constraints are imposed, neural networks might loose a significant portion of their expressive power zhang2022rethinking . Therefore, typically no constraints are imposed, and the neural network’s Lipschitz constant is determined once the model is trained. Unfortunately, even for two-layer neural networks, exact computation of this quantity is NP-hard virmaux2018lipschitz . A recent line of work has thus focused on estimating the Lipschitz constant of neural networks, mainly by deriving upper bounds virmaux2018lipschitz ; fazlyab2019efficient ; latorre2020lipschitz ; combettes2020lipschitz ; kim2021lipschitz ; pabbaraju2021estimating . Efficiency is often sacrificed for the sake of tighter bounds (e. g., use of semidefinite programming) which underlines the need for accurate estimation of the Lipschitz constant.
@@
Prior work estimates mainly the Lipschitz constant of architectures composed of fully-connected and convolutional layers. However, in several domains, input data might correspond to complex objects which consist of other simpler objects. We typically model these complex objects as sets or multisets (i. e., a generalization of a set). For instance, in computer vision, a point cloud is a set of data points in the $3$-dimensional space. Likewise, in natural language processing, documents may be represented by multisets of word embeddings. Neural networks for sets typically consist of a series of fully-connected layers followed by an aggregation function which produces a representation for the entire multiset zaheer2017deep ; qi2017pointnet . For the model to be invariant to permutations of the multiset’s elements, a permutation invariant aggregation function needs to be employed. Common functions include the sum, mean and max operators. While previous work has studied the expressive power of neural networks that employ the aforementioned aggregation functions wagstaff2019limitations ; wagstaff2022universal , their Lipschitz continuity and stability to perturbations remains underexplored. Besides those standard aggregation functions, other methods for embedding multisets have been proposed recently such as the Fourier Sliced-Wasserstein embedding which is bi-Lipschitz with respect to the Wasserstein distance amir2025fourier .
@@
In this paper, we consider three metrics that compute the distance between multisets of vectors and investigate whether the three commonly-employed aggregation functions (i. e., sum, mean, max) are Lipschitz continuous with respect those metrics. We show that for multisets of arbitrary size, each aggregation function is Lipschitz continuous with respect to only a single metric. On the other hand, if all multisets have equal cardinalities, each aggregation function is Lipschitz continuous with respect to other metrics as well. Our results are summarized in Table 1. We also study the Lipschitz constant of neural networks for sets which employ the aforementioned aggregation functions. We find that for multisets of arbitrary size, the models that employ the mean and max operators are Lipschitz continuous with respect to a single metric and we provide upper bounds on their Lipschitz constants. Strikingly, we also find that there exist models that employ the sum operator which are not Lipschitz continuous. We also relate the Lipschitz constant of those networks to their generalization performance under distribution shifts. We verify our theoretical results empirically on real-world datasets from different domains.
@@
Table 1: Summary of main results. Lipschitz constants of the different aggregation functions with respect to the three considered metrics. $d$ denotes the dimension of the vectors. The “–” symbol denotes that the function is not Lipschitz continuous with respect to a given metric. $\dagger$: all multisets have equal cardinalities ($=M$).
@@ 2.1 Notation
[…]
@@ 2.2 Lipschitz Continuous Functions
In this paper, we focus on functions $f\colon\mathcal{S}(\mathbb{R}^{d})\rightarrow\mathbb{R}^{d^{\prime}}$ that map multisets of $d$-dimensional vectors to $d^{\prime}$-dimensional vectors. Therefore, $\mathcal{X}=\mathcal{S}(\mathbb{R}^{d})$, while $\mathcal{Y}=\mathbb{R}^{d^{\prime}}$. For $d_{\mathcal{X}}$, we consider three distance functions for multisets of vectors (presented in subsection 2.4 below), while $d_{\mathcal{Y}}$ is induced by the $\ell_{2}$-norm, i. e., $d_{\mathcal{Y}}\big{(}f(x_{1}),f(x_{2})\big{)}=||f(x_{1})-f(x_{2})||_{2}$.
@@ 2.3 Aggregation Functions
$f_{\text{sum}}(X)=\sum_{\mathbf{v}\in X}\mathbf{v}$ $f_{\text{mean}}(X)=\frac{1}{|X|}\sum_{\mathbf{v}\in X}\mathbf{v}$ $\big{[}f_{\text{max}}(X)\big{]}_{i}=\max\Big{(}\big{\{}[\mathbf{v}]_{i}\colon\mathbf{v}\in X\big{\}}\Big{)},\quad\forall i\in[d]$
It is well-known that sum aggregators can represent injective functions over multisets zaheer2017deep ; xu2019powerful in case the input multisets contain elements from a countable universe. However, it has been empirically observed that mean and max aggregators can outperform the sum aggregator in certain applications zaheer2017deep ; cappart2023combinatorial .
@@ 2.4 Distance Functions for Unordered Multisets
We next present the three considered metrics for comparing multisets to each other. Let $X=\{\!\!\{\mathbf{v}_{1},\ldots,\mathbf{v}_{m}\}\!\!\}$ and $Y=\{\!\!\{\mathbf{u}_{1},\ldots,\mathbf{u}_{n}\}\!\!\}$ denote two multisets of vectors, i. e., $X,Y\in\mathcal{S}(\mathbb{R}^{d})$. The three metrics require to compute the distance between each element of the first multiset and every element of the second multiset. We use the distance induced by the $\ell_{2}$-norm (i. e., Euclidean distance) to that end. Note also that all three metrics can be computed in polynomial time in the number of elements of the input multisets.
@@
The earth mover’s distance (EMD) is a measure of dissimilarity between two distributions. Roughly speaking, given two distributions, the output of EMD is proportional to the minimum amount of work required to change one distribution into the other. Over probability distributions, EMD is also known as the Wasserstein metric with $p=1$ ($\mathcal{W}_{1}$). We use the formulation of the EMD where the total weights of the signatures are equal to each other which is known to be a metric rubner2000earth :
$d_{\text{EMD}}(X,Y)=\min_{\mathbf{F}}$ $\sum_{i=1}^{m}\sum_{j=1}^{n}[\mathbf{F}]_{ij}\,||\mathbf{v}_{i}-\mathbf{u}_{j}||_{2}$
@@
The Hausdorff distance is another measure of dissimilarity between two multisets of vectors. It represents the maximum distance of a multiset to the nearest point in the other multiset, and is defined as follows:
$h(X,Y)=\max_{i\in[m]}\min_{j\in[n]}||\mathbf{v}_{i}-\mathbf{u}_{j}||_{2}$
@@
The $d_{H}(\cdot,\cdot)$ function is a metric, and roughly speaking, its value is small if every point of either set is close to some point of the other set.
@@
where
$M(X,Y)=\min_{\pi\in\mathfrak{S}_{m}}\bigg{[}\sum_{i=1}^{n}||\mathbf{v}_{\pi(i)}-\mathbf{u}_{i}||_{2}+\sum_{i=n+1}^{m}||\mathbf{v}_{\pi(i)}||_{2}\bigg{]}$
and $M(Y,X)$ is defined analogously. If the elements of the input multisets do not contain the zero vector, the matching distance is a metric.
Proposition 2.2 (Proof in Appendix A.1).
The matching distance is a metric on $\mathcal{S}(\mathbb{R}^{d}\setminus\{\mathbf{0}\})$ where $d\in\mathbb{N}$ and $\mathbf{0}$ is the zero vector.
@@
Proposition 2.3 (Proof in Appendix A.2).
Let $X,Y\in\mathcal{S}(\mathbb{R}^{d}\setminus\{\mathbf{0}\})$ denote two multisets of the same size, i. e., $|X|=|Y|=M$. Then, we have that $d_{M}(X,Y)=Md_{\text{EMD}}(X,Y)$.
@@ 3.1 Lipschitz Continuity of Aggregation Functions
Theorem 3.1 (Proof in Appendix A.3).
Let $X,Y\in\mathcal{S}_{\leq M}(\mathbb{R}^{d})$ denote two multisets that consist of $d$-dimensional vectors ($X,Y\in\mathcal{S}_{\leq M}(\mathbb{R}^{d}\setminus\{\mathbf{0}\})$ for matching distance). Then,
@@
    The mean function is Lipschitz continuous with respect to EMD and its Lipschitz constant is $L=1$, but is not Lipschitz continuous with respect to the Hausdorff distance and with respect to the matching distance.
@@
    The sum function is Lipschitz continuous with respect to the matching distance and its Lipschitz constant is $L=1$, but is not Lipschitz continuous with respect to EMD and with respect to the Hausdorff distance.
@@
    The max function is Lipschitz continuous with respect to the Hausdorff distance and its Lipschitz constant is $L=\sqrt{d}$, but is not Lipschitz continuous with respect to EMD and with respect to the matching distance.
The above theoretical result suggests that there is some correspondence between the three aggregation functions and the three metrics for unordered multisets. In fact, each aggregation function seems to be closely related to a single metric. We also observe that while the Lipschitz constants of the sum and mean functions with respect to the matching distance and to EMD, respectively, is constant (equal to $1$), the Lipschitz constant of the max function with respect to the Hausdorff distance depends on the dimension $d$ of the vectors.
As discussed above, the sum function is theoretically more powerful than the rest of the functions. However, in practice it has been observed that in certain tasks mean and max aggregators lead to higher levels of performance zaheer2017deep ; cappart2023combinatorial . While this discrepancy between theory and empirical evidence may be attributed to the training procedure rather than to expressivity, it suggests that there is no single aggregation function that provides a superior performance under all possible circumstances and motivates the study of all three of them.
Our previous result showed that each aggregation function is Lipschitz continuous only with respect to a single metric for multisets. It turns out that if the multisets have fixed size, then the aggregation functions are Lipschitz continuous also with respect to other metrics.
Lemma 3.2 (Proof in Appendix A.4).
Let $\mathcal{X}=\{X_{1},\ldots,X_{N}\}$ denote a set of multisets where $X_{i}\in\mathcal{S}_{M}(\mathbb{R}^{d})$ for all $i\in[N]$ ($X_{i}\in\mathcal{S}_{M}(\mathbb{R}^{d}\setminus\{\mathbf{0}\})$ for matching distance). Then,
@@
    The mean function restricted to inputs from set $\mathcal{X}$ is Lipschitz continuous with respect to the matching distance and its Lipschitz constant is $L=\frac{1}{M}$.
@@
    The sum function restricted to inputs from set $\mathcal{X}$ is Lipschitz continuous with respect to EMD and its Lipschitz constant is $L=M$.
@@
    The max function restricted to inputs from set $\mathcal{X}$ is Lipschitz continuous with respect to EMD and its Lipschitz constant is $L=M$, and it is also Lipschitz continuous with respect to the matching distance and its Lipschitz constant is $L=1$.
According to the above Lemma, the max function is Lipschitz continuous with respect to all three distance metrics when all multisets have the same cardinality. This suggests that in such a setting, if any of the three metrics is insensitive to some perturbation applied to the input data, this perturbation will not result in a large variation in the output of the max function.

### v3

@@ Abstract
The Lipschitz constant of a neural network is connected to several important properties of the network such as its robustness and generalization. It is thus useful in many settings to estimate the Lipschitz constant of a model. Prior work has focused mainly on estimating the Lipschitz constant of multi-layer perceptrons and convolutional neural networks. Here we focus on data modeled as sets or multisets of vectors and on neural networks that can handle such data. These models typically apply some permutation invariant aggregation function, such as the sum, mean or max operator, to the input multisets to produce a single vector for each input sample. In this paper, we investigate whether these aggregation functions, along with an attention-based aggregation function, are Lipschitz continuous with respect to three distance functions for unordered multisets, and we compute their Lipschitz constants. In the general case, we find that each aggregation function is Lipschitz continuous with respect to only one of the three distance functions, while the attention-based function is not Lipschitz continuous with respect to any of them. Then, we build on these results to derive upper bounds on the Lipschitz constant of neural networks that can process multisets of vectors, while we also study their stability to perturbations and generalization under distribution shifts. To empirically verify our theoretical analysis, we conduct a series of experiments on datasets from different domains.
@@ 1 Introduction
In the past decade, deep neural networks have been applied with great success to several problems in different machine learning domains ranging from computer vision (Krizhevsky et al., 2012; He et al., 2016) to natural language processing (Vaswani et al., 2017; Peters et al., 2018). Owing to their recent success, these models are now ubiquitous in machine learning applications. However, deep neural networks can be very sensitive to their input. Indeed, it is well-known that if some specially designed small perturbation is applied to an image, it can cause a neural network model to make a false prediction even though the perturbed image looks “normal” to humans (Szegedy et al., 2014; Goodfellow et al., 2015).
@@
A key metric for quantifying the robustness of neural networks to small perturbations is the Lipschitz constant. Training neural networks with bounded Lipschitz constant has been considered a promising direction for producing models robust to adversarial examples (Tsuzuku et al., 2018; Anil et al., 2019; Trockman and Kolter, 2021). However, if Lipschitz constraints are imposed, neural networks might lose a significant portion of their expressive power (Zhang et al., 2022a). Therefore, typically no constraints are imposed, and the neural network’s Lipschitz constant is determined once the model is trained. Unfortunately, even for two-layer neural networks, exact computation of this quantity is NP-hard (Virmaux and Scaman, 2018). A recent line of work has thus focused on estimating the Lipschitz constant of neural networks, mainly by deriving upper bounds (Virmaux and Scaman, 2018; Fazlyab et al., 2019; Latorre et al., 2020; Combettes and Pesquet, 2020; Kim et al., 2021; Pabbaraju et al., 2021; Chuang and Jegelka, 2022). Efficiency is often sacrificed for the sake of tighter bounds (e. g., use of semidefinite programming) which underlines the need for accurate estimation of the Lipschitz constant.
@@
Prior work estimates mainly the Lipschitz constant of architectures composed of fully-connected and convolutional layers. However, in several domains, input data might correspond to complex objects which consist of other simpler objects. We typically model these complex objects as sets or multisets (i. e., a generalization of a set). For instance, in computer vision, a point cloud is a set of data points in the $3$-dimensional space. Likewise, in natural language processing, documents may be represented by multisets of word embeddings. Neural networks for sets typically consist of a series of fully-connected layers followed by an aggregation function which produces a representation for the entire multiset (Zaheer et al., 2017; Qi et al., 2017a). For the model to be invariant to permutations of the multiset’s elements, a permutation invariant aggregation function needs to be employed. Common functions include the sum, mean and max operators. While previous work has studied the expressive power of neural networks that employ the aforementioned aggregation functions (Wagstaff et al., 2019; 2022), their Lipschitz continuity and stability to perturbations remains underexplored. Besides those standard aggregation functions, other methods for embedding multisets have been proposed recently such as the Fourier Sliced-Wasserstein embedding which is bi-Lipschitz with respect to the Wasserstein distance (Amir and Dym, 2025).
@@
In this paper, we consider three distance functions between multisets of vectors and investigate whether the three commonly-employed aggregation functions (i. e., sum, mean, max), along with an attention-based function, are Lipschitz continuous with respect those functions. We show that for multisets of arbitrary size, each aggregation function is Lipschitz continuous with respect to only a single distance function, while the attention-based function is not Lipschitz continuous with respect to any of them. On the other hand, if all multisets have equal cardinalities, each aggregation function is Lipschitz continuous with respect to other distance functions as well. Our results are summarized in Table 1. We also study the Lipschitz constant of neural networks for sets which employ the aforementioned aggregation functions. We find that for multisets of arbitrary size, the models that employ the mean and max operators are Lipschitz continuous with respect to a single metric and we provide upper bounds on their Lipschitz constants. Strikingly, we also find that there exist models that employ the sum operator which are not Lipschitz continuous with respect to any of the considered functions. We also relate the Lipschitz constant of those networks to their generalization performance under distribution shifts. We verify our theoretical results empirically on real-world datasets from different domains.
@@
Table 1: Summary of main results. Lipschitz constants of the different aggregation functions with respect to the three considered distance functions. $d$ denotes the dimension of the vectors. The “–” symbol denotes that the function is not Lipschitz continuous with respect to a given distance function. $\dagger$: all multisets have equal cardinalities (=M=M).
@@ 2.1 Notation
Let $\mathbb{N}$ denote the set of natural numbers. Then, $n]=\{1,\ldots,n\}\subset\mathbb{N}$ for $n\geq 1$. Let also $\{\!\!\{\}\!\!\}$ denote a multiset, i. e., a generalized concept of a set that allows multiple instances for its elements. Since a set is also a multiset, in what follows we use the term “multiset” to refer to both sets and multisets. Here we focus on finite multisets whose elements are $d$-dimensional real vectors. Let $M\in\mathbb{N}\setminus\{1\}$. We denote by $\mathcal{S}_{\leq M}(\mathbb{R}^{d})$ and by $\mathcal{S}_{M}(\mathbb{R}^{d})$ the set of all those multisets that consist of at most $M$ and of exactly $M$ elements, respectively. We drop the subscript when it is clear from context. The elements of a multiset do not have an inherent ordering. Therefore, the two multisets $X=\{\!\!\{\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{2}\}\!\!\}$ and $Y=\{\!\!\{\mathbf{v}_{2},\mathbf{v}_{1},\mathbf{v}_{2}\}\!\!\}$ are equal to each other, i. e., $X=Y$. The cardinality $|X|$ of a multiset $X$ is equal to the number of elements of $X$. Vectors are denoted by boldface lowercase letters (e. g., $\mathbf{v}$ and $\mathbf{u}$) and matrices by boldface uppercase letters (e. g., $\mathbf{A}$ and $\mathbf{M}$). Given some vector $\mathbf{v}$, we denote by $\mathbf{v}]_{i}$ the $i$-th element of the vector. Likewise, given some matrix $\mathbf{M}$, we denote by $\mathbf{M}]_{ij}$ the element in the $i$-th row and $j$-th column of the matrix.
@@ 2.2 Lipschitz Continuous Functions
In this paper, we focus on functions $f\colon\mathcal{S}(\mathbb{R}^{d})\rightarrow\mathbb{R}^{d^{\prime}}$ that map multisets of $d$-dimensional vectors to $d^{\prime}$-dimensional vectors. Therefore, $\mathcal{X}=\mathcal{S}(\mathbb{R}^{d})$, while $\mathcal{Y}=\mathbb{R}^{d^{\prime}}$. For $d_{\mathcal{X}}$, we consider three distance functions for multisets of vectors (presented in subsection 2.4 below), while $d_{\mathcal{Y}}$ is induced by the $\ell_{2}$-norm, i. e., $d_{\mathcal{Y}}\big(f(x_{1}),f(x_{2})\big)=\|f(x_{1})-f(x_{2})\|_{2}$.
@@ 2.3 Aggregation Functions
$f_{\textsc{sum}}(X)=\sum_{\mathbf{v}\in X}\mathbf{v}$ $f_{\textsc{mean}}(X)=\frac{1}{|X|}\sum_{\mathbf{v}\in X}\mathbf{v}$ $\big[f_{\textsc{max}}(X)\big]_{i}=\max\Big(\big\{[\mathbf{v}]_{i}\colon\mathbf{v}\in X\big\}\Big),\quad\forall i\in[d]$
The sum aggregator can represent a strictly larger class of functions over sets than the mean and max aggregators. If the elements of the input sets come from a countable set $\mathcal{X}$, then for an appropriate $f\colon\mathcal{X}\rightarrow\mathbb{R}$, the function defined as $g(\{x_{1},\ldots,x_{n}\})=\sum_{i=1}^{n}f(x_{i})$ maps the input sets injectively to $\mathbb{R}$ (Zaheer et al., 2017). Notably, it is also shown that injectivity is sufficient for approximation. On the other hand, the mean and max functions are not injective set functions. These results have been also generalized to multisets (Xu et al., 2019). Note, however, that it has been empirically observed that mean and max aggregators can outperform the sum aggregator in certain applications (Zaheer et al., 2017; Cappart et al., 2023).
@@ 2.4 Distance Functions for Unordered Multisets
We next present the three considered functions for comparing multisets to each other. Let $X=\{\!\!\{\mathbf{v}_{1},\ldots,\mathbf{v}_{m}\}\!\!\}$ and $Y=\{\!\!\{\mathbf{u}_{1},\ldots,\mathbf{u}_{n}\}\!\!\}$ denote two multisets of vectors, i. e., $X,Y\in\mathcal{S}(\mathbb{R}^{d})$. The three functions require to compute the distance between each element of the first multiset and every element of the second multiset. We use the distance induced by the $\ell_{2}$-norm (i. e., Euclidean distance) to that end. Note also that all three functions can be computed in polynomial time in the number of elements of the input multisets.
@@
The earth mover’s distance (EMD) is a measure of dissimilarity between two distributions (Rubner et al., 2000). Roughly speaking, given two distributions, the output of EMD is proportional to the minimum amount of work required to change one distribution into the other. Over probability distributions, EMD is also known as the Wasserstein metric with $p=1$ ($\mathcal{W}_{1}$). We use the formulation of the EMD where the total weights of the signatures are equal to each other which is known to be a metric on the space of sets of vectors (Rubner et al., 2000) and a pseudometric on $\mathcal{S}(\mathbb{R}^{d})$:
$d_{\text{EMD}}(X,Y)=\min_{\mathbf{F}}$ $\sum_{i=1}^{m}\sum_{j=1}^{n}[\mathbf{F}]_{ij}\,\|\mathbf{v}_{i}-\mathbf{u}_{j}\|_{2}$
@@
The Hausdorff distance is another measure of dissimilarity between two multisets of vectors (Rockafellar and Wets, 1998). It represents the maximum distance of a multiset to the nearest point in the other multiset, and is defined as follows:
$h(X,Y)=\max_{i\in[m]}\min_{j\in[n]}\|\mathbf{v}_{i}-\mathbf{u}_{j}\|_{2}$
@@
The bidirectional Hausdorff distance is a metric on the space of sets of vectors and a pseudometric on $\mathcal{S}(\mathbb{R}^{d})$. Roughly speaking, its value is small if every point of either set is close to some point of the other set.
@@
$\text{where }\qquad M(X,Y)=\min_{\pi\in\mathfrak{S}_{m}}\bigg[\sum_{i=1}^{n}\|\mathbf{v}_{\pi(i)}-\mathbf{u}_{i}\|_{2}+\sum_{i=n+1}^{m}\|\mathbf{v}_{\pi(i)}\|_{2}\bigg]$
and $M(Y,X)$ is defined analogously. Variants of this distance function have been introduced in prior work (Chuang and Jegelka, 2022; Davidson and Dym, 2024). If the elements of the input multisets do not contain the zero vector, the matching distance is a metric.
Proposition 2.2 (Proof in Appendix B.1).
The matching distance is a metric on $\mathcal{S}(\mathbb{R}^{d}\setminus\{\mathbf{0}\})$ where $d\in\mathbb{N}$ and $\mathbf{0}$ is the zero vector. It is a pseudometric on $\mathcal{S}(\mathbb{R}^{d})$.
@@
Proposition 2.3 (Proof in Appendix B.2).
Let $X,Y\in\mathcal{S}(\mathbb{R}^{d})$ denote two multisets of the same size, i. e., $|X|=|Y|=M$. Then, we have that $d_{M}(X,Y)=Md_{\text{EMD}}(X,Y)$.
@@ 3.1 Lipschitz Continuity of Aggregation Functions
Theorem 3.1 (Proof in Appendix B.3).
@@
    The mean function defined on $\mathcal{S}_{\leq M}(\mathbb{R}^{d})$ is Lipschitz continuous with respect to EMD and its Lipschitz constant is $L=1$, but is not Lipschitz continuous with respect to the Hausdorff distance and with respect to the matching distance.
@@
    The sum function defined on $\mathcal{S}_{\leq M}(\mathbb{R}^{d})$ is Lipschitz continuous with respect to the matching distance and its Lipschitz constant is $L=1$, but is not Lipschitz continuous with respect to EMD and with respect to the Hausdorff distance.
@@
    The max function defined on $\mathcal{S}_{\leq M}(\mathbb{R}^{d})$ is Lipschitz continuous with respect to the Hausdorff distance and its Lipschitz constant is $L=\sqrt{d}$, but is not Lipschitz continuous with respect to EMD and with respect to the matching distance.
The above theoretical result suggests that there is some correspondence between the three aggregation functions and the three distance functions for unordered multisets. In fact, each aggregation function seems to be closely related to a single metric. We also observe that while the Lipschitz constants of the sum and mean functions with respect to the matching distance and to EMD, respectively, is constant (equal to $1$), the Lipschitz constant of the max function with respect to the Hausdorff distance depends on the dimension $d$ of the vectors (which is typically larger than $1$). A direct consequence of the above Theorem is that no two of the considered distance functions are bi-Lipschitz equivalent.
As discussed above, the sum function is theoretically more expressive than the rest of the functions. However, in practice it has been observed that in certain tasks mean and max aggregators lead to higher levels of performance (Zaheer et al., 2017; Cappart et al., 2023). While this discrepancy between theory and empirical evidence may be attributed to the training procedure rather than to expressivity, it suggests that there is no single aggregation function that provides a superior performance under all possible circumstances and motivates the study of all three of them.
Our previous result showed that each aggregation function is Lipschitz continuous only with respect to a single distance function for multisets. It turns out that if the multisets have fixed size, then the aggregation functions are Lipschitz continuous also with respect to other functions. This is not surprising given Proposition 2.3.
Lemma 3.2 (Proof in Appendix B.4).
@@
    The mean function defined on $\mathcal{S}_{M}(\mathbb{R}^{d})$ is Lipschitz continuous with respect to the matching distance and its Lipschitz constant is $L=\frac{1}{M}$, but is not Lipschitz continuous with respect to the Hausdorff distance.
@@
    The sum function defined on $\mathcal{S}_{M}(\mathbb{R}^{d})$ is Lipschitz continuous with respect to EMD and its Lipschitz constant is $L=M$, but is not Lipschitz continuous with respect to the Hausdorff distance.
@@
    The max function defined on $\mathcal{S}_{M}(\mathbb{R}^{d})$ is Lipschitz continuous with respect to EMD and its Lipschitz constant is $L=M$, and it is also Lipschitz continuous with respect to the matching distance and its Lipschitz constant is $L=1$.
According to the above Lemma, the max function is Lipschitz continuous with respect to all three distance functions when all multisets have the same cardinality. This suggests that in such a setting, if any of the three distance functions is insensitive to some perturbation applied to the input data, this perturbation will not result in a large variation in the output of the max function.
In addition to the standard aggregation functions described above, recent work has also explored neural-based approaches for aggregating multisets of vectors. Here, we consider an attention mechanism, which is commonly employed to produce a vector from a multiset of vectors, and has achieved significant success in the fields of natural language processing (Yang et al., 2016; Nikolentzos et al., 2020) and graph learning (Veličković et al., 2018; Brody et al., 2022). Given a multiset $X=\{\!\!\{\mathbf{v}_{1},\ldots,\mathbf{v}_{m}\}\!\!\}\in\mathcal{S}(\mathbb{R}^{d})$, the attention mechanism is defined as follows:
$f_{\textsc{att}}(X)=\sum_{i=1}^{m}\alpha_{i}\mathbf{v}_{i}\quad\text{where}\quad\alpha_{i}$ $=\frac{\exp\big(\mathbf{q}^{\top}g(\mathbf{W}\,\mathbf{v}_{i})\big)}{\sum_{j=1}^{m}\exp\big(\mathbf{q}^{\top}g(\mathbf{W}\,\mathbf{v}_{j})\big)}$
where $\mathbf{W}\in\mathbb{R}^{d^{\prime}\times d}$ and $\mathbf{q}\in\mathbb{R}^{d^{\prime}}$ denote a trainable matrix and a trainable vector, respectively, while $g$ denotes some activation function. The output of the mechanism is a convex combination of the multiset’s elements.
We next investigate whether the attention mechanism $f_{\textsc{att}}(X)$ is Lipschitz continuous with respect to the three considered distance functions for multisets of vectors.
Proposition 3.3 (Proof in Appendix B.5).
There exist instances of $f_{\textsc{att}}(X)$ defined on $\mathcal{S}_{M}(\mathbb{R}^{d})$ which are not Lipschitz continuous with respect to any of the three considered distance functions.
Our result aligns with the finding of Kim et al. (2021) who showed that the standard self-attention mechanism is not Lipschitz. Note, however, that the definition of the considered attention mechanism differs from that of self-attention. Kim et al. (2021) also proposed an alternative $\ell_{2}$ self-attention that is Lipschitz. Incorporating $\ell_{2}$ attention into the definition of $f_{\textsc{att}}(X)$, unfortunately, does not make it Lipschitz (more details in Appendix B.6).